OpenStudy (anonymous):

solve 2x+10=2x medals!!!!!!!!!!!!!!!

OpenStudy (mertsj):

Subtract 2x from both sides.

OpenStudy (anonymous):

so it would be 10=x?

OpenStudy (anonymous):

I've got that one before but it doesn't fit the info i have

OpenStudy (mertsj):

2x-2x=0 NOT x

OpenStudy (anonymous):

Im lost

OpenStudy (anonymous):

how did you go from 2x+10=2x to 2x-2x=0?

OpenStudy (mertsj):

I told you to subtract 2x from both sides. And you said 10=x. So I pointed out that when you subtract 2x-2x on the right side, the result should be 0 not x

OpenStudy (anonymous):

oh ok

OpenStudy (mertsj):